OpenStudy (anonymous):

Let f(x) = -3x - 6 and g(x) = 5x + 2. Find f(x) + g(x) I got 2x-4 but my study partner is saying it's wrong. I'm confused.

jimthompson5910 (jim_thompson5910):

f(x) + g(x) = (-3x - 6) + (5x + 2) f(x) + g(x) = -3x - 6 + 5x + 2 f(x) + g(x) = (-3x + 5x) + (-6 + 2) f(x) + g(x) = (-3 + 5)x + (-6 + 2) f(x) + g(x) = (2)x + (-4) f(x) + g(x) = 2x - 4 you are correct

OpenStudy (anonymous):

Thanks a bunch!

jimthompson5910 (jim_thompson5910):

np